# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 4 The bimodule structure and the Brauer quotients of $\mathbb{A}_{\delta}$
where $X$ is isomorphic to a direct sum of modules of the form $\mbox{$\mathcal{O}$}DhD$ for some $h\in G-N_{G}(D)$. The module $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is independent of the choice of $g$ up to isomorphism and has multiplicity $1$ for any $gDC_{G}(D)\in E_{\mathfrak{h}}$. Moreover, there is a unitary subalgebra $\mathbb{B}$ of $\mathbb{A}_{\delta}$ isomorphic to the twisted group algebra $\mbox{$\mathcal{O}$}_{\alpha}(D\rtimes E_{\mathfrak{h}})$ for some $\alpha$ of $H^{2}(E_{\mathfrak{h}};k^{*})$ such that it contains $\mbox{$\mathcal{O}$}D$ and $\mathbb{A}_{\delta}=\mathbb{B}\oplus Y$ as $\mathbb{B}$-$\mathbb{B}$-bimodules. In particular, $\mathbb{B}$ is isomorphic to $\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ and $Y$ is isomorphic to $X$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules.
@@
By Lemma 4.5 and [29, Lemma (44.7)], we have $\mathbb{A}_{\delta}=(\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g})\oplus X$ such that $\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ and $X$ have no common direct summand. Suppose that $X$ has a direct summand isomorphic $\mbox{$\mathcal{O}$}DhD$ with $h\in N_{G}(D)$. This means that $\mathbb{A}_{\delta}$ has a direct summand $\mbox{$\mathcal{O}$}D\mathfrak{a}_{h}$ such that $u\mathfrak{a}_{h}w^{-1}=\mathfrak{a}_{h}$ if and only if $w=h^{-1}uh$ for any $(u,w)\in D\times D$.
@@
On the other hand, by Lemma 4.1, the element $h$ has to belong to $N_{G}(D_{\delta})$. Then $h=vx$ for some $v\in P$ and $xDC_{G}(D)\in E_{\mathfrak{h}}$. Since there is an element $\mathfrak{a}_{x}$ of $\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{x}(uj)\mathfrak{a}_{x}^{-1}=(xux^{-1}j)$ for any $u\in D$, we can assume that $h=v$ is in $P$. Recall from the proof of Lemma 3.6, for $v\in P$, there is an element $a_{v,\delta}$ of $A_{\delta}^{*}\cap j\mathbb{A}vj$ such that $a_{v,\delta}(uj)a_{v,\delta}^{-1}=(vuv^{-1})j$ for any $u\in D$. Therefore, $\mathfrak{a}_{v}a_{v,\delta}^{-1}$ is in $j\mathbb{A}v^{-1}j$ and $u\mathfrak{a}_{v}a_{v,\delta}^{-1}w^{-1}=\mathfrak{a}_{v}a_{v,\delta}^{-1}$ if and only if $u=w$ for any $(u,w)\in D\times D$. In particular, $\mbox{$\mathcal{O}$}D\mathfrak{a}_{v}a_{v,\delta}^{-1}$ is isomorphic to $\mbox{$\mathcal{O}$}D$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, which implies that $(j\mathbb{A}v^{-1}j)(D)\neq 0$. By Lemma 3.6, we have $v$ is in $DC_{P}(D)$. Then $\mbox{$\mathcal{O}$}Dv$ is isomorphic to $\mbox{$\mathcal{O}$}D$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, from which we can deduce that $X$ has a direct summand isomorphic to $\mbox{$\mathcal{O}$}D\mathfrak{a}_{x}$. This contradicts Lemma 4.5. So the direct summand of $X$ has the form $\mbox{$\mathcal{O}$}DhD$ with $h\in G-N_{G}(D)$. The second statement follows from Lemma 4.5.
Let us consider the remaining statements of this proposition. For any $g\in N_{G}(D_{\delta})$, we denote $gDC_{G}(D)\in N_{G}(D_{\delta})/DC_{G}(D)$ by $\tilde{g}$. By the proof of Proposition 3.7, we have $N_{\mathbb{A}_{\delta}}(D)/D(\mathbb{A}_{\delta}^{D})^{*}$ isomorphic to $E_{\mathfrak{h}}$. In particular, we have the following short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${D(\mathbb{A}_{\delta}^{D})^{*}/(D\times k^{*})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\mathbb{A}_{\delta}}(D)/(D\times k^{*})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
@@
It is clear that $D(\mathbb{A}_{\delta}^{D})^{*}/(D\times k^{*})$ is isomorphic to $j+J(\mathbb{A}_{\delta})^{D}$. Since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group, by [29, Lemma (45.6)], this sequence splits. So there is a subgroup $\hat{E}_{\mathfrak{h}}$ of $N_{\mathbb{A}_{\delta}}(D)$ containing $D\times k^{*}$ such that $\hat{E}_{\mathfrak{h}}\cap(D(\mathbb{A}_{\delta}^{D})^{*})=D\times k^{*}$ and $\hat{E}_{\mathfrak{h}}/(D\times k^{*})\cong E_{\mathfrak{h}}$. This yields another short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(D\times k^{*})/k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\hat{E}_{\mathfrak{h}}/k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
@@
which still splits by the Schur-Zassenhaus Theorem. Now let $\mathbb{M}$ be a subgroup of $\hat{E}_{\mathfrak{h}}$ containing $k^{*}$ such that $\mathbb{M}/k^{*}$ is a complement of $(D\times k^{*})/k^{*}$ in $\hat{E}_{\mathfrak{h}}/k^{*}$. Set $\mathbb{B}$ to be the $\mathcal{O}$-submodule of $\mathbb{A}_{\delta}$ generated by $D\rtimes\mathbb{M}$. It is routine to check that $\mathbb{B}$ is a unitary subalgebra containing $\mbox{$\mathcal{O}$}D$ of $\mathbb{A}_{\delta}$ which is isomorphic to $\mbox{$\mathcal{O}$}_{\alpha}(D\rtimes E_{\mathfrak{h}})$ as $D$-interior algebras for the element $\alpha\in H^{2}(E_{\mathfrak{h}};k^{*})$ determined the second short exact sequence above and isomorphic to $\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Therefore, $\mathrm{Br}_{D}(\mathbb{B})$ is isomorphic to $kZ(D)$. By [21, Lemma 5.5 and Proposition 5.15], there is a finite subgroup $\mathbb{M}^{\prime}$ of $\mathbb{M}$ such that $\mathbb{B}$ is isomorphic to $\mbox{$\mathcal{O}$}(D\rtimes\mathbb{M}^{\prime})\mathfrak{b}^{\prime}$ for some central idempotent $\mathfrak{b}^{\prime}$ of $\mbox{$\mathcal{O}$}(D\rtimes\mathbb{M}^{\prime})$ as $D$-interior algebras. Then $\mathrm{Br}_{D}(\mbox{$\mathcal{O}$}(D\rtimes\mathbb{M}^{\prime})\mathfrak{b}^{\prime})$ is isomorphic to the local $k$-algebra $kZ(D)$ which implies $\mathfrak{b}^{\prime}$ is a block idempotent since $D$ is normal in $D\rtimes\mathbb{M}^{\prime}$. Hence, by [17, Proposition 6.7.15], the inclusion map $\mathbb{B}$ to $\mathbb{A}_{\delta}$ splits as $\mathbb{B}$-$\mathbb{B}$-bimodules. We are done. ∎
By the decomposition (4.6.1), we can directly get the information about the rank of $\mathbb{A}_{\delta}$ and the dimension of simple modules as follows.
@@
Another consequence of Proposition 4.6, we can show that a special automorphism of $\mathbb{A}_{\delta}$ is an inner automorphism. This result is similar to [21, Proposition 14.9]
@@
With the notation above, let $f$ be an $\mathcal{O}$-algebra automorphism of $\mathbb{A}_{\delta}$ such that its restriction to the subalgebra $\mathbb{B}$ is trivial. Then $f$ is an inner automorphism induced by an element of $C_{\mathbb{A}_{\delta}}(\mathbb{B})^{*}$.
@@
Since $\mathbb{A}_{\delta}=\mathbb{B}\oplus Y$ as $\mathbb{B}$-$\mathbb{B}$-bimodules, we have $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ being isomorphic to $\mathbb{A}_{\delta}\oplus Y\oplus(Y\mathop{\otimes}\limits_{\mathbb{B}}Y)$ as $\mathbb{B}$-$\mathbb{B}$-bimodules. Since $Y$ is isomorphic to $X$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, we have $\mathrm{Br}_{D}(Y)\cong\mathrm{Br}_{D}(X)=0$. For $Y\mathop{\otimes}\limits_{\mathbb{B}}Y$, it is clear that it is isomorphic to a direct summand of $Y\mathop{\otimes}\limits_{\mathcal{O}D}Y$ as $\mathbb{B}$-$\mathbb{B}$-bimodules and then it is isomorphic to $X\mathop{\otimes}\limits_{\mathcal{O}D}X$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. By the same argument in the proof of [10, Lemma 3.3], we have $\mathrm{Br}_{D}(X\mathop{\otimes}\limits_{\mathcal{O}D}X)=0$. So is $\mathrm{Br}_{D}(Y\mathop{\otimes}\limits_{\mathbb{B}}Y)$. Since $\mathbb{B}$ contains $\mbox{$\mathcal{O}$}D$ and $\mathbb{A}_{\delta}$ is relatively $\mbox{$\mathcal{O}$}D$-separable, it is easy to check that $\mathbb{A}_{\delta}$ is also relatively $\mathbb{B}$-separable. In particular, $\mathbb{A}_{\delta}$ is isomorphic to a direct summand of $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. Then $\mathbb{A}_{\delta}$ is the unique direct summand of the $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodule $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$, up to isomorphisms, with respect to the Brauer quotients at $D$ being nonzero.
@@
On the other hand, since $\mathbb{A}_{\delta}$ is relatively $\mathbb{B}$-separable, then the multiplication map $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}\longrightarrow\mathbb{A}_{\delta}$ splits. This means that there is an element $\sum\limits_{l}\mathfrak{a}_{l}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\mathfrak{a}^{\prime}_{l}$ of $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ such that $\sum\limits_{l}\mathfrak{a}_{l}\mathfrak{a}^{\prime}_{l}=j$ and $\sum\limits_{l}\mathfrak{a}\mathfrak{a}_{l}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\mathfrak{a}^{\prime}_{l}=\sum\limits_{l}\mathfrak{a}_{l}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\mathfrak{a}^{\prime}_{l}\mathfrak{a}$ for any $\mathfrak{a}\in\mathbb{A}_{\delta}$. Let $(\mathbb{A}_{\delta})_{f}$ be the $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodule such that it equals $\mathbb{A}_{\delta}$ as left $\mathbb{A}_{\delta}$-modules with $f(\mathfrak{a})$ acting by right multiplication on $\mathfrak{a}^{\prime}\in(\mathbb{A}_{\delta})_{f}$ as $\mathfrak{a}$ on $\mathfrak{a}^{\prime}\in\mathbb{A}_{\delta}$. With the same arguments in the proof of [10, Proposition 3.11], $(\mathbb{A}_{\delta})_{f}$ is isomorphic to a direct summand of $(\mathbb{A}_{\delta})_{f}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. Since the restriction of $f$ to $\mathbb{B}$ is trivial, it is easy to check that $(\mathbb{A}_{\delta})_{f}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ is isomorphic to $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. Obviously, $(\mathbb{A}_{\delta})_{f}$ is isomorphic to $\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. By the uniqueness of $\mathbb{A}_{\delta}$, $(\mathbb{A}_{\delta})_{f}$ is also isomorphic to $\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. The image of $j$ under this isomorphism is the element we want. ∎

### v2

@@ 4 The bimodule structure and the Brauer quotients of $\mathbb{A}_{\delta}$
where $X$ is isomorphic to a direct sum of $\mbox{$\mathcal{O}$}(D\times D)$-modules of the form $\mbox{$\mathcal{O}$}DhD$ for some $h\in G-N_{G}(D)$. The $\mbox{$\mathcal{O}$}(D\times D)$-module $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is independent of the choice of $g$ up to isomorphism and has multiplicity $1$ for any $gDC_{G}(D)\in E_{\mathfrak{h}}$. Moreover, there is a unitary subalgebra $\mathbb{B}$ of $\mathbb{A}_{\delta}$ isomorphic to the twisted group algebra $\mbox{$\mathcal{O}$}_{\alpha}(D\rtimes E_{\mathfrak{h}})$ for some $\alpha$ of $Z^{2}(E_{\mathfrak{h}};k^{\times})$ such that it contains $\mbox{$\mathcal{O}$}Dj$ and $\mathbb{A}_{\delta}=\mathbb{B}\oplus Y$ as $\mathbb{B}$-$\mathbb{B}$-bimodules. In particular, $\mathbb{B}$ and $Y$ are isomorphic to $\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ and $X$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, respectively.
@@
By Lemma 4.5 and [29, Lemma (44.7)], we have $\mathbb{A}_{\delta}=(\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g})\oplus X$ such that $\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ and $X$ have no common direct summand up to isomorphism. Suppose that $X$ has a direct summand isomorphic to $\mbox{$\mathcal{O}$}DhD$ with $h\in N_{G}(D)$. This implies that there is an element $\mathfrak{a}_{h}$ in $\mathbb{A}_{\delta}$ satisfying that $u\mathfrak{a}_{h}w^{-1}=\mathfrak{a}_{h}$ if and only if $w=h^{-1}uh$ for any $(u,w)\in D\times D$ and the $\mbox{$\mathcal{O}$}(D\times D)$-module $\mbox{$\mathcal{O}$}D\mathfrak{a}_{h}$ is a direct summand of $\mathbb{A}_{\delta}$.
@@
On the other hand, by Lemma 4.1, the element $h$ has to belong to $N_{G}(D_{\delta})$. Then $h=vx$ for some $v\in P$ and $xDC_{G}(D)\in E_{\mathfrak{h}}$. Since there is an element $\mathfrak{a}_{x}$ in $\mathbb{A}_{\delta}^{\times}$ such that $\mathfrak{a}_{x}(uj)\mathfrak{a}_{x}^{-1}=(xux^{-1}j)$ for any $u\in D$, $\mbox{$\mathcal{O}$}D\mathfrak{a}_{h}\mathfrak{a}_{x}^{-1}$ is also a direct summand of $\mathbb{A}_{\delta}$ which is isomorphic to $\mbox{$\mathcal{O}$}Dv$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Recall from the proof of Lemma 3.6, for $v\in P$, there is an element $a_{v,\delta}$ of $A_{\delta}^{\times}\cap j\mathbb{A}vj$ such that $a_{v,\delta}(uj)a_{v,\delta}^{-1}=(vuv^{-1})j$ for any $u\in D$. Therefore, $\mathbb{A}_{\delta}a_{v,\delta}^{-1}$ is contained in $j\mathbb{A}v^{-1}j$ and has a direct summand isomorphic to $\mbox{$\mathcal{O}$}D$. In particular, it implies that $(j\mathbb{A}v^{-1}j)(D)\neq 0$. By Lemma 3.6, we have $v$ is in $DC_{P}(D)$. So we can get that $X$ has a direct summand isomorphic to $\mbox{$\mathcal{O}$}D\mathfrak{a}_{x}$, which is impossible. So the direct summand of $X$ has the form $\mbox{$\mathcal{O}$}DhD$ with $h\in G-N_{G}(D)$. The second statement follows from Lemma 4.5.
@@
Let us consider the remaining statements of this proposition. Since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group, we can identify it with a subgroup of $N_{G}(D_{\delta})/C_{G}(D)$. For any $g\in N_{G}(D_{\delta})$, we denote $gC_{G}(D)\in N_{G}(D_{\delta})/C_{G}(D)$ by $\tilde{g}$. Recall from the paragraph 4.4 that for any $\tilde{g}\in E_{\mathfrak{h}}$, there is an element $\mathfrak{a}_{g}$ in $\mathbb{A}_{\delta}^{\times}$ such that $\mathfrak{a}_{g}(uj)\mathfrak{a}_{g}^{-1}=(gug^{-1})j$ for any $u\in D$. Denote the subset $\{\mathfrak{a}_{g}\mathfrak{a}_{\delta}\mid\tilde{g}\in E_{\mathfrak{h}},\mathfrak{a}_{\delta}\in(\mathbb{A}_{\delta}^{D})^{\times}\}$ by $\hat{E}_{\mathfrak{h}}$. Clearly, $\hat{E}_{\mathfrak{h}}$ is a subgroup of $N_{\mathbb{A}_{\delta}}(D)$ containing $(\mathbb{A}_{\delta}^{D})^{\times}$ such that $\hat{E}_{\mathfrak{h}}/(\mathbb{A}_{\delta}^{D})^{\times}$ is isomorphic to $E_{\mathfrak{h}}$. Since $(\mathbb{A}_{\delta}^{D})^{\times}$ is isomorphic to $(j+J(\mathbb{A}_{\delta}^{D}))\times k^{\times}$, we can get the following short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${j+J(\mathbb{A}_{\delta}^{D})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\hat{E}_{\mathfrak{h}}/k^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
By [29, Lemma (45.6)], this sequence splits since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group. Now let $\mathbb{M}$ be a subgroup of $\hat{E}_{\mathfrak{h}}$ containing $k^{\times}$ such that $\mathbb{M}/k^{\times}$ is a complement of $j+J(\mathbb{A}_{\delta}^{D})$ in $\hat{E}_{\mathfrak{h}}/k^{\times}$. This yields another short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{M}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
@@
which determines an element $\alpha$ in $Z^{2}(E_{\mathfrak{h}};k^{\times})$. Set $\mathbb{B}$ to be the $\mathcal{O}$-submodule of $\mathbb{A}_{\delta}$ generated by $Dj\rtimes\mathbb{M}$. It is routine to check that $\mathbb{B}$ is a unitary subalgebra containing $\mbox{$\mathcal{O}$}Dj$ of $\mathbb{A}_{\delta}$ which is isomorphic to $\mbox{$\mathcal{O}$}_{\alpha}(D\rtimes E_{\mathfrak{h}})$ as $D$-interior algebras and isomorphic to $\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Therefore, $\mathrm{Br}_{D}(\mathbb{B}^{D})$ is isomorphic to $kZ(D)$. By [21, Lemma 5.5 and Proposition 5.15], there is a finite subgroup $\mathbb{M}^{\prime}$ of $\mathbb{M}$ such that $\mathbb{B}$ is isomorphic to $\mbox{$\mathcal{O}$}(Dj\rtimes\mathbb{M}^{\prime})\mathfrak{b}^{\prime}$ for some central idempotent $\mathfrak{b}^{\prime}$ of $\mbox{$\mathcal{O}$}(Dj\rtimes\mathbb{M}^{\prime})$ as $D$-interior algebras. Then $\mathrm{Br}_{D}((\mbox{$\mathcal{O}$}(Dj\rtimes\mathbb{M}^{\prime})\mathfrak{b}^{\prime})^{D})$ is isomorphic to the local $k$-algebra $kZ(D)$ which implies $\mathfrak{b}^{\prime}$ is a block idempotent since $Dj$ is normal in $Dj\rtimes\mathbb{M}^{\prime}$. Hence, by [16, Proposition 6.7.15], the inclusion map $\mathbb{B}$ to $\mathbb{A}_{\delta}$ is split injective as a homomorphism of $\mathbb{B}$-$\mathbb{B}$-bimodules. We are done. ∎
Before giving a refinement of this $\mbox{$\mathcal{O}$}(D\times D)$-module structure for the hyperfocal subalgebra of a hyperfocal abelian Frobenius block, we state three consequences of Proposition 4.6. First, we can get the information about the rank of $\mathbb{A}_{\delta}$ and the dimension of simple modules as follows.
@@
As the second consequence of Proposition 4.6, we can show that a special automorphism of $\mathbb{A}_{\delta}$ is an inner automorphism. This result is similar to [21, Proposition 14.9] and may be of independent interest.
@@
With the notation above, let $f$ be an $\mathcal{O}$-algebra automorphism of $\mathbb{A}_{\delta}$ such that its restriction to the subalgebra $\mathbb{B}$ is trivial. Then $f$ is an inner automorphism induced by an element in $C_{\mathbb{A}_{\delta}}(\mathbb{B})^{\times}$. Here, $C_{\mathbb{A}_{\delta}}(\mathbb{B})^{\times}$ denotes the centralizer of $\mathbb{B}$ in $\mathbb{A}_{\delta}^{\times}$.
@@
Since $\mathbb{A}_{\delta}=\mathbb{B}\oplus Y$ as $\mathbb{B}$-$\mathbb{B}$-bimodules, we have $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ being isomorphic to $\mathbb{A}_{\delta}\oplus Y\oplus(Y\mathop{\otimes}\limits_{\mathbb{B}}Y)$ as $\mathbb{B}$-$\mathbb{B}$-bimodules. Since $Y$ is isomorphic to $X$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, we have $\mathrm{Br}_{D}(Y)\cong\mathrm{Br}_{D}(X)=0$. For $Y\mathop{\otimes}\limits_{\mathbb{B}}Y$, it is clear that it is isomorphic to a direct summand of $Y\mathop{\otimes}\limits_{\mathcal{O}D}Y$ as $\mathbb{B}$-$\mathbb{B}$-bimodules and then it is isomorphic to $X\mathop{\otimes}\limits_{\mathcal{O}D}X$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. By the same argument in the proof of [10, Lemma 3.3], we have $(X\mathop{\otimes}\limits_{\mathcal{O}D}X)(D)=0$. So is $(Y\mathop{\otimes}\limits_{\mathbb{B}}Y)(D)$. Since $\mathbb{B}$ contains $\mbox{$\mathcal{O}$}D$ and $\mathbb{A}_{\delta}$ is relatively $\mbox{$\mathcal{O}$}D$-separable, it is easy to check that $\mathbb{A}_{\delta}$ is also relatively $\mathbb{B}$-separable. In particular, $\mathbb{A}_{\delta}$ is isomorphic to a direct summand of $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. Then $\mathbb{A}_{\delta}$ is the unique direct summand, up to isomorphism, of the $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodule $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$, with respect to the Brauer quotients at $D$ being nonzero.
@@
On the other hand, since $\mathbb{A}_{\delta}$ is relatively $\mathbb{B}$-separable, then the multiplication map $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}\longrightarrow\mathbb{A}_{\delta}$ splits. This means that there is an element $\sum\limits_{l}\mathfrak{a}_{l}\mathop{\otimes}\limits_{\mathbb{B}}\mathfrak{a}^{\prime}_{l}$ in $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ such that $\sum\limits_{l}\mathfrak{a}_{l}\mathfrak{a}^{\prime}_{l}=j$ and $\sum\limits_{l}\mathfrak{a}\mathfrak{a}_{l}\mathop{\otimes}\limits_{\mathbb{B}}\mathfrak{a}^{\prime}_{l}=\sum\limits_{l}\mathfrak{a}_{l}\mathop{\otimes}\limits_{\mathbb{B}}\mathfrak{a}^{\prime}_{l}\mathfrak{a}$ for any $\mathfrak{a}\in\mathbb{A}_{\delta}$. Let $(\mathbb{A}_{\delta})_{f}$ be the $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodule such that it equals $\mathbb{A}_{\delta}$ as left $\mathbb{A}_{\delta}$-modules with $f(\mathfrak{a})$ acting by right multiplication on $\mathfrak{a}^{\prime}\in(\mathbb{A}_{\delta})_{f}$ as $\mathfrak{a}$ on $\mathfrak{a}^{\prime}\in\mathbb{A}_{\delta}$. With the same arguments in the proof of [10, Proposition 3.11], $(\mathbb{A}_{\delta})_{f}$ is isomorphic to a direct summand of $(\mathbb{A}_{\delta})_{f}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. Since the restriction of $f$ to $\mathbb{B}$ is trivial, it is easy to check that $(\mathbb{A}_{\delta})_{f}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ is isomorphic to $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. Obviously, $(\mathbb{A}_{\delta})_{f}$ is isomorphic to $\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. By the uniqueness of $\mathbb{A}_{\delta}$, $(\mathbb{A}_{\delta})_{f}$ is also isomorphic to $\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. The image of $j$ under this isomorphism is the element we want. ∎
The third consequence is about the equivariant property of $\mathbb{A}$-modules. To state this property, we need some notation for twisted modules.